OpenStudy (anonymous):

Find the number of terms in the following arithmetic sequence. 1, 3, 5, 7, . . ., 51

OpenStudy (anonymous):

Looks like a(n)=1+2(n-1), n=1, a(1)=1 n=2, a(2)=3 so n=26, a(26)=1+2(26-1)=51 Therefore 26 terms in the sequence.

OpenStudy (anonymous):

this is an example of an arithmetic progression, where every adjacent values have a common difference... the general formula is given by... a(n) = a(1) + (n - 1) x d where: a(1) = the value of the 1st term, a(n) = the value of the nth term, n = the number of terms in the given sequence, d = the common difference between adjacent values in a sequence....

OpenStudy (anonymous):

from the solution of @DominicNg... you just plug-in the given values... and you can solve for n....